An employee is paid a salary of $11,050 per month. If the current FICA rate is 6.2% on the first $130,000 of earnings, and the Medicare tax rate is 1.45% of all earnings, how much should be withheld for (a) FICA tax and (b) Medicare tax during the following months? March and December.
Answer #1

a.

March:

FICA tax should be withheld for the whole month’s salary, since the limit $130,000 is not yet reached.

FICA tax = 11,050 × 6.2%

                = 685.10 (Answer)

December:

March to December is a gap of 10 months; total earnings is (11,050 × 10 =) 110,500; it is still below the maximum limit of $130,000. Therefore,

FICA tax = 11,050 × 6.2%

                = 685.10 (Answer)

b.

This would be on total earnings, without any limit restriction.

March:

Medicare tax = 11,050 × 1.45%

                        = 160.225

                        = 160.23 (Answer)

December:

Medicare tax = 11,050 × 1.45%

                        = 160.225

                        = 160.23 (Answer)